Let f(x)={x^2-16x, x<9;12rootx, x> or equal to 9.Is f continuous at x=9? Determine whether f is differentiable at x=9.If so, find the value of the derivative there.
1
Expert's answer
2020-05-11T18:33:49-0400

1.

"f(x) = \begin{cases}
x^2 - 16x & \text{x<9} \\
12 \sqrt{x} & \text{x $\ge 9$}
\end{cases}" . To determine whether f is continuous at x=9, we should check if: "\lim_{x \to 9-0} f(x) = \lim_{x \to 9+0} f(x) = f(9)" , or in our terms: (since f differs when x<9 and x"\ge"9 ):

"\lim_{x \to 9-0} x^2 -16x = \lim_{x \to 9+0} 12 \sqrt{x} = f(9)."

Since "f(x) = 12 \sqrt{x} , x \ge 9," and "12 \sqrt{x}" is continuous at every "x \in R" , we obtain that "\lim_{x \to 9+0} 12 \sqrt{x} = f(9) = 36." So, we should check whether "\lim_{x \to 9-0} x^2 - 16x = 36" . We will use the fact that "x^2 - 16x" is continuous at every "x \in R" , and thus at x=9: "\lim_{x->9-0}x^2 -16x = \lim_{x \to 9+0} x^2 -16x = 9^2 - 16 \cdot 9 = 81 - 144 = -63" . So, we obtain that "\lim_{x \to 9-0} f(x) \neq \lim_{x \to 9+0} f(x)". Hence, f is not continuous at x=9.


2.

Let's assume f is differentiable at x=9. We will use the fact that if f is differentiable at point "x_0" then f is continuous at "x_0" .

Hence, if f is differentiable at x=9, f should be contionuous at x=9, but we proved it is not. Thus, f can not be differentiable at x=9.